# Imaging heat transport in suspended diamond nanostructures with integrated spin defect thermometers

### v1

@@ I.1 Sample preparation
We then define the cantilevers, aligned to the Cr patches, using a second round of electron beam lithography followed by several etching steps. A thin titanium layer (40 nm) is evaporated on the surface. A 2 µm-thick layer of HSQ (FOx-16) is spin coated above and exposed with electron beam lithography followed by the aforementioned development process. The Ti layer serves as a conductive and adhesive layer between the diamond the HSQ. A Chlorine-based reactive ion etching process allows to pattern the Ti layer (STS Multiplex ICP, 800 W ICP power, 150 W bias power, 10 sccm Cl2 and 10 sccm BCl3, 3 mTorr). To transfer the pattern from the HSQ hard mask to the diamond, a vertical etch is applied with directional oxygen plasma (STS Multiplex ICP, 400 W ICP power, 200 W bias power, 50 sccm O2, 15 mTorr). Faraday cage angled etch is then used under the same plasma condition as the previous step to undercut the structures [65]. The hard mask is stripped in the HF 1% bath.
Since the Faraday cage angled etch technique restricts the lateral dimensions of suspended structures, we proceed further with focused ion beam (FIB) milling. This method allows greater flexibility in structuring, enabling us to precisely polish narrow cantilevers and undercut wider ones [66, 67]. A 200 nm Ti layer is deposited to shield the sample from gallium (Ga) implantation and to prevent drift during milling caused by charge accumulation. The milling is conducted with a Zeiss CrossBeam 540 SEM/FIB dual beams system using 30 kV Ga+ ions on both sides of each cantilever. We apply cross-section mode with 1.5 nA beam current to sputter away most of the material and lower the ion current to 700 pA to realize better sidewall smoothness. Finally, cantilevers with finely polished sidewalls are revealed after removing the Ti layer in HF bath, as shown in Fig. 1(b).
@@ I.2 Experimental setup
All experiments are carried out in our home-built confocal microscope. The diamond sample is mounted on a temperature-controlled stage, kept at a constant temperature of 22 for experiments in Fig. 2 and Fig. 4. The output beam of a 515nm laser is split into two beams, one for heating the cantilever and the other for readout of the NV centers’ ZFS. The power of each beam can be controlled independently. The two beams are focused on different positions of the cantilevers with a high numerical aperture microscope objective (NA=0.9). PL of the NV centers is collected into a multi-mode fiber and detected with a multi-pixel photon counter (Hamamatsu C14455-1550GA). Microwave are delivered to the sample with a wire loop placed in close vicinity to the diamond sample. Microwave frequency is modulated directly within the microwave generator (Rhodes&Schwarz SMB 100A), with sinusoidal modulation shape, to use in combination with the lock-in amplifier. The modulation period is 0.2 ms.
The readout parameters are chosen to optimize temperature sensitivity: the readout laser power is 0.2 mW, microwave power 41 dBm and microwave frequency modulation depth 5 MHz (corresponding roughly to half the ODMR linewidth). We have checked that the above parameters do not affect the readout ZFS value, confirming that the readout does not induce local heating of the cantilever. We have also verified that the heating laser does not induce any change in ZFS when focused on other portions of the cantilever, away from the metal patch.
@@ I.3 Independent study of the patch temperature
To confirm that the metal patch is well thermalized to the diamond cantilever, we measure the hot carrier anti-Stokes emission of the patch to get an independent estimate of its temperature. Luminescence in noble metals following photoexcitation is a complex process influenced by various mechanisms, including electron-electron scattering, electron-phonon coupling, multiphoton excitations, and surface plasmons[68, 69]. Anti-Stokes emission corresponds to photons emitted at a higher energy than the energy of the laser incident on the material. As detailed in previous studies, the anti-Stokes emission spectrum can be fit with Boltzmann statistics to determine the effective temperature [68, 70]. The spectrum is given by:
where and represent emission probability and a density of states respectively and represents the anti-Stokes temperature. The exponential term, or Boltzmann factor, merges the energy contribution from electrons and phonons. The term varies much slower than the Boltzmann factor [71, 70], hence Eq. (S1) can be simplified as:
@@
We measure the anti-Stokes emission spectrum on three different cantilevers, using 532nm continuous wave excitation. This experiment was performed on a different sample than discussed in the main text, where the metal patch was made of gold (Au) instead of Cr. Still, a few nm-thin Cr layer was used as adhesion layer between Au and diamond. Hence, aside from the absorption coefficient of the patch , this difference is not expected to lead to any major qualitative difference in local heat generation and patch thermalization. Anti-stokes emission is measured on a spectrometer. A notch filter rejects most of the reflected excitation laser. Collected spectra are shown in Fig. S1 (a) and can be well fitted with the exponential decay from Eq. (S2) over the range that we can experimentally access. Note that we discard the energy window corresponding to the Raman peak of diamond, to avoid spurious residual contribution from the diamond cantilever. We extract the anti-Stokes temperature from the fit.
@@
The relationship between and the lattice temperature is in general not straightforward, specifically under continuous wave illumination [72]. Hence, we perform a calibration measurement to relate the extracted to the patch temperature, using the temperature-controlled stage on which the diamond sample is mounted (see inset of Fig. S1(b)). We find that the relation between and the controlled patch temperature can be well described using a linear fit: . We can therefore determine the temperature of gold patches versus heating laser power, on cantilevers with different widths, as shown in Fig. S1(b). Although this method does not guarantee high precision (note also that the extracted zero-laser-power temperature is significantly lower on the 1.18 µm cantilever), this measurement proves that the metal patch temperature increases linearly when heating laser is applied. Temperature estimates with anti-Stokes emission are also found to be on the same order as values presented in the main text on diamond cantilevers for similar heating power, meaning that the interface thermal resistance between the patch and diamond is small. Consequently, convective and radiative thermal losses in the patch can safely be neglected.
I.4 Estimation of experimental effective thermal conductivity
@@
The effective thermal conductivity of our cantilevers is defined as . The temperature gradient is directly extracted from the measured temperature profile. We must only estimate the heat flux density in order to obtain . As introduced in the main text, we have . The heating laser power incident on the Cr patch is measured directly. The Cr patch absorption coefficient is estimated by measuring its reflection coefficient and assuming negligible transmission and scattering, i.e. . Finally, the cantilever cross-section is , where both the width and cross-section angle are estimated from SEM images of the different cantilevers, see for example Fig. 3(a) of the main text.
@@
Note that we use an uncertainty of 50 nm in the extracted width from SEM images. This uncertainty needs to be properly taken into account in the uncertainty of , together with the uncertainty on the temperature gradient. We find that the width uncertainty is the dominant factor, explaining the significantly bigger error bars in , see Fig. 4, compared to the temperature gradient in Fig. 3(c).
@@ I.6 Theoretical analysis
where is phonon distribution (, with being the phonon wavevector and the phonon mode index). The LBTE is a microscopic model, which describes the propagation and scattering of phonons in the presence of a temperature gradient. The LBTE can be parametrized from first principles, and its linear-response solution allows to predict the thermal conductivity. We note, in passing, that the LBTE has been recently generalized to account not only for the ability of phonons to propagate particle-like, but also also for their capability to tunnel wave-like [76, 77, 78]. However, in diamond the LBTE is sufficient to accurately describe thermal transport, since in this case propagation dominates over tunneling [12].
The first-principles solution of the LBTE has the drawback of being computationally expensive, making it unpractical to explore how shape effects affect heat transport. Therefore, to investigate thermal transport in cantilevers having different shapes and sizes, we rely on the mesoscopic viscous heat equations (VHE) [12]. The VHE are two coupled partial differential equations for the mesoscopic temperature () and phonon drift-velocity () fields. They are obtained from a coarse-graining the LBTE[79, 12]. The VTE encompass both the diffusive (Fourier) and hydrodynamic regimes [80, 79] of thermal transport, as well as the crossover between them. It has been shown that the VHE rationalizes the experimental demonstration of hydrodynamic transport in graphite [12], and also yields predictions compatible with the space-dependent solution of the LBTE[52, 53] in complex-shaped devices [54]. In the steady-state regime, for a material with isotropic transport properties such as diamond, the VHE read:
@@
where is the the average heat flux discussed in the main text.
@@
The importance on viscuous effects can thus be estimated by comparing the results of computing the heat flux with Eq. (S5) (full solution) versus Eq. (S8) (purely diffusive case).
@@
Finally, we discuss how we considered the Ga contamination in the vicinity of milled faces of the cantilevers (i.e. bottom faces, see inset of Fig. 4). Since with the methods currently available it is not possible to solve the LBTE with position-dependent Ga content, we approximatively consider uniform Ga impurities over the whole cantilever cross-section. This approximation allows to capture the increased influence of the Ga-contaminated layer (of fixed depth , see main text) when reducing the cantilever width . In particular, we consider an impurity concentration proportional to the ratio of contaminated area to total cross-section area, referred to as penetration ratio
@@
In summary, we determine the linear-response solution of the LBTE (Eq. (S3)) for diamond nanowires with width-dependent Ga impurity scattering using the ShengBTE package [55] and use the bulk thermal viscosity computed following the theoretical prescriptions of Ref. [12]. Then, we solve the VHE numerically (Eq. (S4)) using a finite-element solver implemented in Mathematica [82], imposing a temperature gradient K around room temperature on µm long cantilevers. Finally, we obtain the effective thermal conductivity from Eqs (S5) and (S6). The Ga concentration in contaminated areas is set in the ShengBTE simulation as to fit the experimental trend of vs width (black squares in Fig. 4 of the main text). This is the sole fitting parameter in our model.
@@ I.7 Evaluation of the Ga participation ratio
Given that Ga impurities are mostly localized on the two slanted sides of Diamond cantilevers (see inset of Fig. 4), we can compute the participation ratio, PR, as the ratio between the azure area in the triangle on the right of Fig. 4, , and the total triangular section, . Considering an angle at the surface equal to , with some algebra one finds

### v2

@@ I.1 Sample preparation
We then define the cantilevers, aligned to the Cr patches, using a second round of electron beam lithography followed by several etching steps. A thin titanium layer (40 nm) is evaporated on the surface. A 2 µm-thick layer of HSQ (FOx-16) is spin coated above and exposed with electron beam lithography followed by the aforementioned development process. The Ti layer serves as a conductive and adhesive layer between diamond and HSQ. A Chlorine-based reactive ion etching process is used to pattern the Ti layer (STS Multiplex ICP, 800 W ICP power, 150 W bias power, 10 sccm Cl2 and 10 sccm BCl3, 3 mTorr). To transfer the pattern from the HSQ hard mask to the diamond, a vertical etch is applied with directional oxygen plasma (STS Multiplex ICP, 400 W ICP power, 200 W bias power, 50 sccm O2, 15 mTorr). Faraday cage angled etch is then used under the same plasma condition as the previous step to undercut the structures [70]. The hard mask is stripped in the HF 1% bath.
Since the Faraday cage angled etch technique restricts the lateral dimensions of suspended structures, we proceed further with focused ion beam (FIB) milling. This method allows greater flexibility in structuring, enabling us to precisely polish narrow cantilevers and undercut wider ones [71, 72]. A 200 nm Ti layer is deposited to shield the sample from gallium (Ga) implantation and to prevent drift during milling caused by charge accumulation. The milling is conducted with a Zeiss CrossBeam 540 SEM/FIB dual beams system using 30 kV Ga+ ions on both sides of each cantilever. We apply cross-section mode with 1.5 nA beam current to sputter away most of the material and lower the ion current to 700 pA to realize better sidewall smoothness. Finally, cantilevers with finely polished sidewalls are revealed after removing the Ti layer in HF bath, as shown in Fig. 1b.
@@ I.2 Experimental setup
All experiments are carried out in our home-built confocal microscope. The diamond sample is mounted on a temperature-controlled stage, kept at a constant temperature of 22 ∘C for experiments in Fig. 2 and Fig. 3. The output beam of a 515 nm laser is split into two beams, one for heating the chromium patch at the apex of the cantilever and the other one for reading out the NV centers’ fluorescence all along the cantilever. The power of each beam can be controlled independently. The two beams are focused on different positions of the cantilevers with a high numerical aperture microscope objective (NA=0.9). The fluorescence is collected into a multi-mode fiber and detected with a multi-pixel photon counter (Hamamatsu C14455-1550GA).
Microwaves are delivered to the sample with a wire loop placed in close vicinity to the diamond sample. The microwave frequency is sinusoidally modulated at 5 kHz by the microwave generator (Rhodes&Schwarz SMB 100A) to be used in combination with the lock-in amplifier. The readout parameters are chosen to optimize temperature sensitivity: the readout laser power is 0.2 mW, the microwave power is 41 dBm and the microwave frequency modulation depth is 5 MHz (corresponding to approximately half of the ODMR linewidth). We have checked that the above parameters do not affect the measured value of the zero-field splitting (ZFS), confirming that the spin readout induces negligible local heating inside the cantilever. We have also verified that the heating laser does not induce any change in ZFS when focused on other portions of the cantilever, away from the metal patch.
@@ I.3 Independent study of the patch temperature
To confirm that the metal patch is well thermalized with the diamond cantilever, we measure the hot carrier anti-Stokes emission of the patch to get an independent estimate of its temperature. Luminescence in noble metals following photoexcitation is a complex process influenced by various mechanisms, including electron-electron scattering, electron-phonon coupling, multiphoton excitations, and surface plasmons [73, 74]. Anti-Stokes emission corresponds to photons emitted at a higher energy than the energy of the laser incident on the material. As detailed in previous studies, the anti-Stokes emission spectrum can be fit with Boltzmann statistics to obtain a first approximation of the effective temperature [73, 75]. Under the assumption that both the photonic and electronic densities of states vary much slower than the Boltzmann factor close to the laser energy [76, 75], we have the simple expression for the anti-Stokes intensity as a function of the Raman shift
@@
We measure the anti-Stokes emission spectrum on three different cantilevers, using 532 nm continuous wave excitation and a spectrometer. A notch filter rejects most of the reflected excitation laser. Collected spectra are shown in Fig. S1a and can be well fitted with the exponential decay from Eq. (S1) over the range that we can experimentally access. Note that we discard the energy window corresponding to the Raman peak of diamond, to avoid spurious residual contribution from the diamond cantilever. We extract the anti-Stokes temperature from the fit. This experiment was performed on a patch made of gold (Au) instead of Cr, but with a few nm-thin Cr layer used as adhesion layer between Au and diamond. Hence, aside from the absorption coefficient of the patch , no major qualitative difference is expected in the metal-to-diamond thermalization efficiency.
@@
The relationship between and the lattice temperature is in general not straightforward, specifically under continuous wave illumination [77]. Hence, we perform a calibration measurement to relate the extracted to the patch temperature, using the temperature-controlled stage on which the diamond sample is mounted (see inset of Fig. S1b). We find that the relation between and the controlled patch temperature can be well described using a linear fit: . We can therefore determine the temperature of metal patches versus heating laser power, on cantilevers with different widths, as shown in Fig. S1b. Although this method does not grant high precision (note also that the extracted zero-laser-power temperature is significantly lower on the 1.18 µm cantilever), this measurement proves that the metal patch temperature increases linearly when heating laser is applied. Temperature estimates with anti-Stokes emission are also found to be of the same order as values presented in the main text on diamond cantilevers for similar heating power, meaning that the interface thermal resistance between the patch and diamond is sufficiently small. Our main conclusion is that convective and radiative thermal losses from the metal patch to the surrounding can safely be neglected.
I.4 Experimental estimate of the effective thermal conductivity
@@
The effective thermal conductivity of our cantilevers is defined as . The temperature gradient is directly extracted from the measured temperature profile. We only need to estimate the heat flux density in order to obtain . As introduced in the main text, we have . The heating laser power incident on the Cr patch is measured directly. The Cr patch absorption coefficient is estimated by measuring its reflection coefficient and assuming negligible transmission and scattering, i.e. . Finally, the cantilever cross-section is , where both the width and cross-section angle are estimated from SEM images of the different cantilevers, see for example Fig. 3a of the main text.
@@
We assume an uncertainty of 50 nm in the extracted width from SEM images. This uncertainty needs to be properly taken into account in the uncertainty of , together with the uncertainty on the temperature gradient. We find that the width uncertainty is the dominant factor, explaining the significantly bigger error bars in , see Fig. 4, compared to the temperature gradient in Fig. 3c.
@@ I.6 Theoretical analysis
where is phonon distribution (, with being the phonon wavevector and the phonon mode index). The LBTE is a microscopic model, which describes the propagation and scattering of phonons in the presence of a temperature gradient. The LBTE can be parametrized from first principles, and its linear-response solution allows to predict the thermal conductivity. We note, in passing, that the LBTE has been recently generalized to account not only for the ability of phonons to propagate particle-like, but also for their capability to tunnel wave-like [81, 82, 83]. In diamond, propagation dominates over tunneling and the LBTE accurately describes thermal transport [14].
The first-principles solution of the LBTE has the drawback of being computationally expensive, making it unpractical to explore how nanostructure shape affects heat transport. Therefore, to investigate thermal transport in cantilevers having different cross-section shapes and sizes, we rely on the mesoscopic viscous heat equations (VHE) [14]. The VHE are two coupled partial differential equations for the mesoscopic temperature () and phonon drift-velocity () fields. They are obtained from coarse-graining the LBTE[84, 14]. The VTE encompass both the diffusive (Fourier) and hydrodynamic regimes [7, 84] of thermal transport, as well as the crossover between them. It has been shown that the VHE rationalizes the experimental demonstration of hydrodynamic transport in graphite [14], and also yields predictions compatible with the space-dependent solution of the LBTE[57, 58] in complex-shaped devices [59]. In the steady-state regime, for a material with isotropic transport properties such as diamond, the VHE read:
@@
where is the average heat flux discussed in the main text.
@@
The importance of viscous effects can thus be estimated by comparing the results of computing the heat flux with Eq. (S4) (full solution) versus Eq. (S7) (purely diffusive case).
@@
Finally, we discuss how we considered the Ga contamination in the vicinity of milled faces of the cantilevers (i.e. bottom faces, see inset of Fig. 4). Since with the methods currently available it is not possible to solve the LBTE with position-dependent Ga content, we consider uniform Ga impurities over the whole cantilever cross-section. This approximation allows to capture the increased relative influence of the Ga-contaminated layer (of fixed depth , see main text) when reducing the cantilever width . In particular, we consider an impurity concentration proportional to the ratio of contaminated area to total cross-section area, referred to as penetration ratio
@@
In summary, we determine the linear-response solution of the LBTE (Eq. (S2)) for diamond nanowires with width-dependent Ga impurity scattering using the ShengBTE package [60] and use the bulk thermal viscosity computed following the theoretical prescriptions of Ref. [14]. Then, we solve the VHE numerically (Eq. (S3)) using a finite-element solver implemented in Mathematica [86], imposing a temperature gradient K around room temperature on µm long cantilevers. Finally, we obtain the effective thermal conductivity from Eqs (S4) and (S5). The Ga concentration in the contaminated areas is set in the ShengBTE simulation so as to fit the experimental trend of vs width (black squares in Fig. 4 of the main text). This is the sole fitting parameter in our model.
@@ I.7 Evaluation of the Ga participation ratio
Given that Ga impurities are mostly localized on the two slanted sides of Diamond cantilevers (see inset of Fig. 4), we can compute the participation ratio, PR, as the ratio between the azure area in the triangle on the right of Fig. 4, , to the total triangular section, . Considering an angle at the surface equal to , with some algebra one finds